5. Use Greens function to solve -() cos(), y(0)y(1) 0
Answer #1


given boundary value poro blom 9n addi Paga boundasy conditions ylo)=y0)-0-12) Og Ca (2) with the Satisfy Considesr the assoc 9n addition to the above poropesty (e), the pro posed Satisfy Grean s Sunchor mustals o the dollowing thoree proposches CiDThen we knownthat the Solution of the given boundaoy value poroblon c)-a) Y(oc) Glxt) plt)dt () where plx) = y So that lt)= GP= $70, t= 10 y2ans, n = 6/0 119 +Sinc p2afPtdt Sino tan) 2SintGtdt 2 2 (Cospc-tanl Sinec P 4241S 2- +Sinc rut G12t)dt +Sinzcis an iode (h) dantty Eauating to 3aro tha C2 74is -Cost 1lence C=Sint C201Cyt th b-91 Sint by (13) b2-92=-Cot by(13) 17) Sol